OpenStudy (anonymous):

Find all solutions to the equation (sinx+cosx)(sinx– cosx)=3.

OpenStudy (anonymous):

sin^2-cos^2x=3 (1-cos^2x)-cos^2x=3 1-2cos^2x=3 -2cos^2x=2 cos^2x=-1 no solution

OpenStudy (anonymous):

What do you mean no solution?

OpenStudy (anonymous):

sin^2-cos^2x=3 (1-cos^2x)-cos^2x=3 1-2cos^2x=3 -2cos^2x=2 cos^2x=-1 no solution (refer 0)

OpenStudy (anonymous):

Because cos(x) can't be an imaginary number. It must be between -1 and +1.

OpenStudy (anonymous):

Alternate method, maybe less painful.\[3=(sinx+cosx)(sinx– cosx)=\sin^2x-\cos^2x \le \sin^2x+\cos^2x=1\]Contradiction (3 is not less than one) implies no solution. Nothing wrong with methods above, just thought I would put a slightly different method into the discussion.